# Cert-SSB: Toward Certified Sample-Specific Backdoor Defense

### v1

@@ IV-C2 Robust Training Process
Once the optimized noise $\sigma_{\bm{x}}^{*}$ is obtained, we incorporate it into the training process to enhance robustness. Specifically, we first sample $M$ sets of noise vectors $b_{1},\cdots,b_{M}$ from the distribution $D\sim{\textstyle\prod_{i=1}^{n}}\mathcal{N}(0,I)$, where each set contains $n=|\mathcal{D}|$ i.i.d. vectors corresponding to the size of the training dataset. For each sampled noise set $b_{m}$, we construct a perturbed (poisoned) training dataset $\mathcal{D}_{p}+\sigma_{\bm{x}}^{*}b_{m}$ by introducing the sample-specific noise $\sigma_{\bm{x}}^{*}$. Next, we train $M$ smoothed models on these perturbed datasets, denoted as $g_{1}(\bm{x},\mathcal{D},\sigma_{\bm{x}}^{*}),\ldots,g_{M}(\bm{x},\mathcal{D},\sigma_{\bm{x}}^{*})$. To maintain consistency between the noise distributions during training and inference, we further sample a noise vector $\mu_{M}\sim\mathcal{N}(0,{\sigma_{0}}^{2}I_{h})$ for each model $g_{m}$, using the hash value of its trained model file as the random seed. This noise vector is stored alongside the model parameters and added to the input during inference. By introducing noise in both the training and testing phases, we ensure that the ensemble of smoothed models $\{g_{1},\ldots,g_{M}\}$ avoids performance degradation when classifying clean inputs. See Algorithm 1 in our Appendix -A for training details.
IV-D Cert-SSB Inference: Storage-update-based Certification
@@
In this stage, we aggregate the ensemble of smoothed classifiers trained in the first stage to form the final prediction. However, under the use of sample-specific noise $\sigma_{\bm{x}}^{*}$, traditional certification methods are no longer applicable. To address this, we propose a novel storage-update-based certification method. By introducing a ‘storage’ mechanism, this method dynamically adjusts certification regions to ensure they are non-overlapping across inputs while preserving prediction consistency for each individual sample.
Formally, given a trained model $g(\bm{x},\mathcal{D},\sigma_{\bm{x}}^{*})$ and a testing input $\bm{x}$, the prediction is computed via majority voting with the optimized noise. Specifically, we estimate the class probability as $\mathcal{P}_{\bm{\epsilon}}(g(\bm{x},\mathcal{D},\sigma_{\bm{x}}^{*})=y)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\{g_{m}(\bm{x}+\mu_{m},\mathcal{D}_{p}+\sigma_{\bm{x}}^{*}b_{m})=y\}$, where $\mu_{m}$ is the deterministic noise vector sampled during training, and $\mathbb{I}\{\cdot\}$ is the indicator function. The following theorem states the robustness guarantee:
Theorem 1 (Certified Robustness of Cert-SSB).
@@
Let $\mathcal{B}_{\bm{x}}\in R^{d}$ and let $\bm{\delta}:=(\bm{\Delta}_{1},\bm{\Delta}_{2},...,\bm{\Delta}_{n})$ for backdoor patterns $\bm{\Delta}_{i}\in R^{d}$, and let $\mathcal{D}$ be a training set, and let smoothing noise $\hat{Z}\sim\mathcal{N}(0,I)$, $\hat{D}\sim\mathcal{N}(0,I)$. Let $y_{A}\in\mathcal{Y}$, such as $y_{A}=g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}+\bm{\delta})$ with class probabilities satisfying $\mathcal{P}_{\bm{\epsilon}(\hat{Z},\hat{D})}(f(\bm{x}+\mathcal{B}_{\bm{x}}+\sigma_{\bm{x}}^{*}\hat{Z},\mathcal{D}+\bm{\delta}+\sigma_{\bm{x}}^{*}\hat{D})=y_{A})\geq P_{A}\geq P_{B}\geq\max_{y\neq y_{A}}\mathcal{P}_{\bm{\epsilon}(\hat{Z},\hat{D})}(f(\bm{x}+\mathcal{B}_{\bm{x}}+\sigma_{\bm{x}}^{*}\hat{Z},\mathcal{D}+\bm{\delta}+\sigma_{\bm{x}}^{*}\hat{D})=y)$. Then, we have $g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D})=g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}+\bm{\delta})=y_{A}$ for all backdoor patterns $\bm{\Delta}$ satisfying $\sqrt{\sum_{i=1}^{n}\left\|\bm{\Delta}_{i}\right\|_{2}^{2}}\leq r(g;\sigma_{\bm{x}_{i}}^{*})$, where the certified robust radius $r$ is given by
$\begin{array}[]{ll}r(g;\sigma_{\bm{x}_{i}}^{*})=\frac{\sqrt{\sum_{i=1}^{n}(\sigma_{\bm{x}_{i}}^{*})^{2}}}{2}\left(\Phi^{-1}(P_{A}(\sigma_{\bm{x}_{i}}^{*}))-\Phi^{-1}(P_{B}(\sigma_{\bm{x}_{i}}^{*}))\right).\end{array}$ (11)
@@
Notably, traditional certification methods become inapplicable in our setting due to two core limitations: (1) they assume non-overlapping certification regions for all inputs, and (2) they require a globally fixed robustness parameter ($e.g.$, noise level $\sigma$). To overcome these limitations, we propose a storage-update-based certification strategy. Before introducing the proposed method, we first provide a brief definition of the concepts of ‘overlapping’ and ‘non-overlapping’ certification regions. Based on this definition, we introduce Definition 3, which classifies the possible types of certification region overlapping. Then, following this classification, we describe the storage-update mechanism in detail (see Proposition 1).
@@
Let $g$ be a sample-specific smoothed classifier, and let $r(\sigma_{\bm{x}_{1}}^{*})$ denote the certification radius of $g$ at input $\bm{x}_{1}$. For any other input $\bm{x}_{2}$, if $\|\bm{x}_{1}-\bm{x}_{2}\|_{2}\leq r(\sigma_{\bm{x}_{1}}^{*})$, the certification regions of $\bm{x}_{1}$ and $\bm{x}_{2}$ are said to overlapping; otherwise, they are said to be non-overlapping.
Figure 5: Storage-update-based certification (see proposition 1). Given two inputs $\bm{x}_{1}$ and $\bm{x}_{2}$ with certified regions $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$, respectively, there are three cases based on Definition 3.
Definition 3 (Classification Criteria of Certification Regions).
Let the triplet storage set $\mathcal{S}=\{(\bm{x}_{i},\mathcal{Y}_{i},\mathcal{R}_{i})\}_{i=1}^{n}$ store all previously predicted inputs $\bm{x}_{i}$, their corresponding predictions $\mathcal{Y}_{i}$, and their associated certification regions $\mathcal{R}_{i}$. Here, $\mathcal{R}_{i}$ denotes the certification region centered at $\bm{x}_{i}$, characterized by the certification radius $r_{i}$. The certification regions $\mathcal{R}_{i}$ for different inputs are classified as follows (see Figure 5):
  • Case 1: Non-overlapping Certification Regions. All certification regions are non-overlapping, $i.e.$, $\forall i\neq j,\mathcal{R}_{i}\cap\mathcal{R}_{j}=\emptyset$, and the corresponding predictions are different, $i.e.$, $\mathcal{Y}_{i}\neq\mathcal{Y}_{j}$.
  • Case 2: Overlapping Certification Regions with Consistent Predictions. The certification region $\mathcal{R}_{n+1}$ of a new input $\bm{x}_{n+1}$ overlaps with an existing certification region $\mathcal{R}_{i}$, and their predictions are consistent, $i.e.$, $\exists i$ such that $\mathcal{R}_{i}\cap\mathcal{R}_{n+1}\neq\emptyset$ and $\mathcal{Y}_{n+1}=\mathcal{Y}_{i}$.
  • Case 3: Overlapping Certification Regions with Inconsistent Predictions. The certification region $\mathcal{R}_{n+1}$ of a new input $\bm{x}_{n+1}$ overlaps with an existing region $\mathcal{R}_{i}$, but their predictions differ, $i.e.$, $\mathcal{Y}_{n+1}\neq\mathcal{Y}_{i}$. This case can be further divided into two subcases:
    • The new input lies inside the existing certification region, $i.e.$, $\bm{x}_{n+1}\in\mathcal{R}_{i}$ and $\mathcal{R}_{n+1}\cap\mathcal{R}_{i}\neq\emptyset$.
    • The new input lies outside the existing certification region, $i.e.$, $\bm{x}_{n+1}\notin\mathcal{R}_{i}$ but $\mathcal{R}_{n+1}\cap\mathcal{R}_{i}\neq\emptyset$.
TABLE I: Certified performance ($i.e.$, ERA and AER) is reported using the best results across different noise levels at each radius under the all-to-one setting. We evaluate three types of attacks (one-pixel, four-pixel, and blending) on the MNIST, CIFAR-10, and ImageNette datasets. In particular, we mark the best certification results in boldface.
@@
Cert-SSB 0.64 94.94 83.46 58.66 46.30 34.52 20.22 5.76 0
TABLE II: Certified performance ($i.e.$, CRA and ACR) is reported using the best results across different noise levels at each radius under the all-to-one setting. We evaluate three types of attacks (one-pixel, four-pixel, and blending) on the MNIST, CIFAR-10, and ImageNette datasets. In particular, we mark the best certification results in boldface.
@@
Cert-SSB 0.48 49.10 42.86 38.96 34.48 27.62 19.06 6.74 0
Based on the classification in Definition 3, we propose a storage-update-based certification method, that enforces non-overlapping certification regions while maintaining prediction consistency ($i.e.$, $\forall i\neq j,\mathcal{R}_{i}\cap\mathcal{R}_{j}=\emptyset,\mathcal{Y}_{i}\neq\mathcal{Y}_{j}$). In this way, the certification regions of inputs with different predicted labels do not overlapping within the storage set $\mathcal{S}$. This is a key property of a sound certification process.
Now, we introduce Proposition 1, which formalizes the proposed storage-update-based certification method.
Proposition 1 (Storage-update-based Certification).
Based on Definition 3, the storage-update-based certification method handles new inputs according to the following cases:
  • Case 1: If $\forall i\neq j$, $\mathcal{R}_{i}\cap\mathcal{R}_{j}=\emptyset$ and $\mathcal{Y}_{i}\neq\mathcal{Y}_{j}$, then all existing triplets $(\bm{x}_{i},\mathcal{Y}_{i},\mathcal{R}_{i})$ and $(\bm{x}_{j},\mathcal{Y}_{j},\mathcal{R}_{j})$ in storage remain unchanged.
  • Case 2: If there exists some $i$ such that $\mathcal{R}_{n+1}\cap\mathcal{R}_{i}\neq\emptyset$ and $\mathcal{Y}_{n+1}=\mathcal{Y}_{i}$, then the new certification triplet $(\bm{x}_{n+1},\mathcal{Y}_{n+1},\mathcal{R}_{n+1})$ can be directly added to the storage.
  • Case 3: If $\mathcal{R}_{n+1}\cap\mathcal{R}_{i}\neq\emptyset$ and $\mathcal{Y}_{n+1}\neq\mathcal{Y}_{i}$, the method proceeds as follows [^1] This process is straightforward when the certification regions are $\ell_{2}$-balls. (see Figure 5):
    • If $\bm{x}_{n+1}\in\mathcal{R}_{i}$: The new certification region is updated to the largest subset ~$\tilde{\mathcal{R}}_{n+1}$ such that ~$\tilde{\mathcal{R}}_{n+1}\subseteq\mathcal{R}_{n+1}$ and ~$\tilde{\mathcal{R}}_{n+1}\subseteq\mathcal{R}_{i}$. Then, $\mathcal{R}_{n+1}$ is replaced by ~$\tilde{\mathcal{R}}_{n+1}$, and the label $\mathcal{Y}_{n+1}$ is updated to $\mathcal{Y}_{i}$ to ensure prediction consistency.
    • If $\bm{x}_{n+1}\notin\mathcal{R}_{i}$: The new certification region is updated to the largest subset $\mathcal{R}_{n+1}^{\prime}$ such that $\mathcal{R}_{n+1}^{\prime}\subseteq\mathcal{R}_{n+1}$ and $\mathcal{R}_{n+1}^{\prime}\cap\mathcal{R}_{i}=\emptyset$. Then, original $\mathcal{R}_{n+1}$ is replaced by $\mathcal{R}_{n+1}^{\prime}$.
After applying the appropriate case, the final triplet $(\bm{x}_{n+1},\mathcal{Y}_{n+1},\mathcal{R}_{n+1})$ (or its updated form) is added to the storage set $\mathcal{S}$ for use in future certification.
In practice, in our experiments, we did not observe any cases where inputs with different predictions have overlapping certified regions. That is, for each input, the certified region stored in $\mathcal{S}$ is essentially determined by the certification radius computed using Eq. (11) for the sample-specific smoothed classifier $g(\bm{x},\mathcal{D},\sigma_{\bm{x}}^{*})$. This can be attributed to two main reasons: 1) Due to the high dimensionality of image datasets, the $\ell_{2}$-norm distance between samples is significantly larger than the certification radius provided by randomized smoothing; 2) The optimized noise $\sigma_{\bm{x}}^{*}$ tends to have a moderate value ($\sigma_{\bm{x}}^{*}\leq 1.0$), resulting in relatively small certification regions. For example, the certification region corresponds to an $\ell_{2}$-ball with a radius of approximately $4\sigma_{\bm{x}}^{*}$, which is much smaller than the distances between samples in high-dimensional datasets ($e.g.$, ImageNet). Nonetheless, in certain rare or specifically constructed scenarios, it remains theoretically possible for inputs with different predictions to exhibit overlapping certified regions, particularly in low-dimensional input spaces or when abnormally large values of $\sigma_{\bm{x}}^{*}$ lead to significantly enlarged certification radii. For such cases, our storage-update-based certification strategy becomes indispensable: by appropriately adjusting the certified regions and associated predictions, it effectively resolves potential conflicts, as described in Proposition 1.

### v2

@@ IV-C2 Robust Training Process
Once the optimized sample-specific noise $\{\sigma_{\bm{x}_{i}}^{*}\}_{i=1}^{n}$ is obtained, we incorporate it into the training process to enhance robustness. Specifically, we first sample $M$ sets of noise vectors $b_{1},\cdots,b_{M}$ from the distribution $D\sim{\textstyle\prod_{i=1}^{n}}\mathcal{N}(0,I)$, where each set contains $n=|\mathcal{D}|$ i.i.d. vectors corresponding to the size of the training dataset. For each sampled noise set $b_{m}$, we construct a perturbed (poisoned) training dataset $\mathcal{D}_{p}^{(m)}\triangleq\mathcal{D}_{p}+\{\sigma_{\bm{x}_{i}}^{*}b_{m,i}\}_{i=1}^{n}$ by perturbing each data point in $\mathcal{D}_{p}$ with its optimized noise scale $\sigma_{\bm{x}_{i}}^{*}$. Here, $\mathcal{D}_{p}$ denotes the poisoned training dataset consisting of both poisoned and benign samples, as defined in Section III-A. Next, we train $M$ smoothed models on these perturbed datasets, denoted as $g_{1}(\bm{x},\mathcal{D}_{p}^{(1)},\sigma_{\bm{x}}^{*}),\ldots,g_{M}(\bm{x},\mathcal{D}_{p}^{(M)},\sigma_{\bm{x}}^{*})$. To maintain consistency between the noise distributions used during training and inference, for each trained model $g_{m}$, we deterministically sample and store a unit-scale base noise vector $\mu_{m}\sim\mathcal{N}(0,I_{d})$ using a random seed derived from a hash value of the trained model parameters, where $I_{d}$ denotes the $d$-dimensional identity matrix corresponding to the input space ($i.e.$, $d$ is the dimensionality of the input feature vector). This base noise vector is stored together with the model parameters and reused during inference. The detailed inference-time noise application will be described in Section IV-D. By introducing noise perturbations during both training and inference, we ensure that the ensemble of smoothed models $\{g_{1},\ldots,g_{M}\}$ avoids performance degradation when classifying clean inputs. See Algorithm 1 in our Appendix -A for training details.
IV-D Cert-SSBD Inference: Storage-update-based Certification
@@
Building upon the ensemble of smoothed models trained in Section IV-C, we now present the inference and certification procedure of Cert-SSBD. At inference stage, we aggregate the ensemble outputs under the optimized, sample-specific noise scale $\sigma_{\bm{x}}^{*}$ to obtain the final prediction. Since certification is no longer performed with a single fixed noise parameter, existing certification methods are not directly applicable. To address this, we propose a novel storage-update-based certification method. By introducing a ‘storage’ mechanism, this method dynamically adjusts certification regions to ensure they are non-overlapping across inputs while preserving prediction consistency for each individual sample.
Formally, given trained models $\{(g_{m},\mu_{m})\}_{m=1}^{M}$ and a testing input $\bm{x}_{i}$, the prediction is obtained via majority voting under the optimized, sample-specific noise scale $\{\sigma_{\bm{x}_{i}}^{*}\}_{i=1}^{n}$. Concretely, for each model $g_{m}$, we evaluate the prediction on the perturbed input $\bm{x}+\sigma_{\bm{x}}^{*}\mu_{m}$, where $\mu_{m}\sim\mathcal{N}(0,I_{d})$ is the unit-scale base noise vector deterministically sampled and stored during training, and the model is associated with the corresponding perturbed poisoned training dataset $\mathcal{D}_{p}^{(m)}$ (defined in Section IV-C). The resulting vote frequency over the ensemble serves as an unbiased empirical estimate of the class probabilities under the smoothed classifier: $\mathcal{P}_{\bm{\epsilon}}\!\left(g(\bm{x},\{\mathcal{D}_{p}^{(m)}\}_{m=1}^{M},\sigma_{\bm{x}}^{*})=y\right)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\!\left\{g_{m}\!\left(\bm{x}+\sigma_{\bm{x}}^{*}\mu_{m},\;\mathcal{D}_{p}^{(m)}\right)=y\right\}$, where $\mathbb{I}\{\cdot\}$ denotes the indicator function. To account for the statistical uncertainty induced by the finite ensemble size $M$, we estimate one-sided $(1-\alpha)$-binomial confidence bounds on the class probability estimates. Specifically, let $cnts[y]$ denote the number of votes received by class $y$, and let $(y_{A},y_{B})$ be the two classes with the largest vote counts. Based on $cnts[y_{A}]$ and $cnts[y_{B}]$, we compute a lower confidence bound $P_{A}$ for the target class $y_{A}$ and an upper confidence bound $P_{B}$ for the runner-up class at confidence level $\alpha$. If $P_{A}>P_{B}$, the prediction $y_{A}$ is certified with confidence $1-\alpha$, which further enables the computation of a certified robust radius. The following theorem establishes the robustness guarantee.
Figure 5: Storage-update-based certification. Given two inputs $\bm{x}_{1}$ and $\bm{x}_{2}$ with certified regions $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$, and predicted labels $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$, respectively, the certification process may fall into the three configurations illustrated in the figure, as described in Remark 1. Case 1: the storage set remains unchanged. Case 2: the new input together with its certified region can be directly added to the storage set. Case 3: the process is further divided into two sub-cases, (a) $\bm{x}_{2}\in\mathcal{R}_{1}$ (inside) and (b) $\bm{x}_{2}\notin\mathcal{R}_{1}$ (outside). In this case, the certified region of the new input is shrunk to a subset (either restricted to the overlapping part or to the largest non-overlapping subset), so that the certification remains valid while conflicts in the storage set are removed.
Theorem 1 (Certified Robustness of Cert-SSBD).
@@
Let $\mathcal{B}_{\bm{x}}\in\mathbb{R}^{d}$ denote a backdoor trigger applied to the test input, and $\bm{\delta}\triangleq(\bm{\Delta}_{1},\bm{\Delta}_{2},\ldots,\bm{\Delta}_{n})$ denote the collection of training-set perturbations, where $\bm{\Delta}_{i}\in\mathbb{R}^{d}$ and $\bm{\Delta}_{i}=\bm{0}$ for benign training samples (as defined in Section III-A). Let $\mathcal{D}$ be a training set and let smoothing noise $\hat{Z}\sim\mathcal{N}(0,I)$, $\hat{D}\sim\mathcal{N}(0,I)$. Let $y_{A}\in\mathcal{Y}$, such as $y_{A}=g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}+\bm{\delta})$ with class probabilities satisfying $\mathcal{P}_{\bm{\epsilon}(\hat{Z},\hat{D})}(f(\bm{x}+\mathcal{B}_{\bm{x}}+\sigma_{\bm{x}}^{*}\hat{Z},\mathcal{D}+\bm{\delta}+\sigma_{\bm{x}}^{*}\hat{D})=y_{A})\geq P_{A}\geq P_{B}\geq\max_{y_{B}\neq y_{A}}\mathcal{P}_{\bm{\epsilon}(\hat{Z},\hat{D})}(f(\bm{x}+\mathcal{B}_{\bm{x}}+\sigma_{\bm{x}}^{*}\hat{Z},\mathcal{D}+\bm{\delta}+\sigma_{\bm{x}}^{*}\hat{D})=y)$. Then, we have $g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D})=g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}+\bm{\delta})=y_{A}$ for all training-set perturbations $\bm{\delta}$ satisfying $\sqrt{\sum_{i=1}^{n}\left\|\bm{\Delta}_{i}\right\|_{2}^{2}}\leq r(g;\sigma_{\bm{x}}^{*})$, where the certified robust radius $r$ is given by
$\begin{array}[]{ll}r(g;\sigma_{\bm{x}}^{*})=\frac{\sigma_{\bm{x}}^{*}}{2}\left(\Phi^{-1}(P_{A}(\sigma_{\bm{x}}^{*}))-\Phi^{-1}(P_{B}(\sigma_{\bm{x}}^{*}))\right).\end{array}$ (11)
@@
Notably, unlike prior randomized smoothing–based certification methods that rely on a fixed noise level ($i.e.$, a single initialized $\sigma_{0}$ shared by all inputs), Cert-SSBD performs inference under sample-specific optimized noise scales $\sigma_{\bm{x}}^{*}$. Under this setting, the certified region associated with each input is determined by its own noise scale rather than a unified initialized parameter, and therefore the certified regions of different inputs are no longer guaranteed to be globally non-overlapping. When certified regions corresponding to different predicted labels overlap, this may introduce ambiguities in the certification process. Motivated by this understanding, we introduce a storage-update-based certification method to ensure the reliability and soundness of certification under sample-specific noise. To formalize the above issue, we first rigorously define the notions of “overlapping” and “non-overlapping” certification regions.
@@
Let $g$ be a sample-specific smoothed classifier, and let $r(\sigma_{\bm{x}_{1}}^{*})$ denote the certification radius of $g$ at input $\bm{x}_{1}$. For any other input $\bm{x}_{2}$, if $\|\bm{x}_{1}-\bm{x}_{2}\|_{2}\leq r(\sigma_{\bm{x}_{1}}^{*})$, the certification regions of $\bm{x}_{1}$ and $\bm{x}_{2}$ are said to be overlapping; otherwise, they are said to be non-overlapping.
Remark 1.
In general, under sample-specific noise, certifying a new input may lead to three possible configurations of certification regions (see Figure 5): Case 1 (non-overlapping certification regions), Case 2 (overlapping certification regions with consistent predictions), and Case 3 (overlapping certification regions with inconsistent predictions). Among them, Case 3 may lead to ambiguity in certification and therefore requires explicit conflict resolution to ensure soundness.
TABLE I: Certified performance ($i.e.$, ERA and AER) of Cert-SSBD and RAB on MNIST, CIFAR-10, and ImageNette under the all-to-one setting with representative attacks ($i.e.$, one-pixel, four-pixel, and blending). At each radius, we report the best result over different noise levels; the best results are marked in boldface.
@@
Cert-SSBD 0.64 94.94 83.46 58.66 46.30 34.52 20.22 5.76 0
TABLE II: Certified performance ($i.e.$, CRA and ACR) of Cert-SSBD and RAB on MNIST, CIFAR-10, and ImageNette under the all-to-one setting with representative attacks ($i.e.$, one-pixel, four-pixel, and blending). At each radius, we report the best result over different noise levels; the best results are marked in boldface.
@@
Cert-SSBD 0.48 49.10 42.86 38.96 34.48 27.62 19.06 6.74 0
To address the potential overlap of certification regions defined above, we introduce a storage-update-based certification strategy, which enforces non-overlapping certification regions across inputs with different predicted labels while maintaining prediction consistency. Specifically, we maintain a triplet storage set $\mathcal{S}=\{(\bm{x}_{i},\mathcal{Y}_{i},\mathcal{R}_{i})\}_{i=1}^{n}$, where each triplet records a previously certified input $\bm{x}_{i}$, its predicted label $\mathcal{Y}_{i}$, and its associated certification region $\mathcal{R}_{i}$. The strategy requires the storage set to satisfy the following key property: for any $i\neq j$, whenever $\mathcal{Y}_{i}\neq\mathcal{Y}_{j}$, it must hold that $\mathcal{R}_{i}\cap\mathcal{R}_{j}=\emptyset$. This property is necessary to ensure the soundness of the certification process. In particular, given a newly certified triplet $(\bm{x}_{n+1},\mathcal{Y}_{n+1},\mathcal{R}_{n+1})$, if there exists a stored certification region $\mathcal{R}_{i}$ such that $\mathcal{R}_{n+1}\cap\mathcal{R}_{i}\neq\emptyset$ and $\mathcal{Y}_{n+1}\neq\mathcal{Y}_{i}$ (corresponding to Case 3 in Figure 5), we resolve the conflict according to the following two cases: if $\bm{x}_{n+1}\in\mathcal{R}_{i}$, we update the prediction at $\bm{x}_{n+1}$ to $\mathcal{Y}_{i}$ and refine $\mathcal{R}_{n+1}$ to the largest subset consistent with $\mathcal{R}_{i}$; otherwise, if $\bm{x}_{n+1}\notin\mathcal{R}_{i}$, we refine $\mathcal{R}_{n+1}$ to the largest subset that does not intersect with $\mathcal{R}_{i}$. This procedure is applied sequentially over all elements in the storage set, after which the updated triplet is added to $\mathcal{S}$. As a result, certification regions associated with different predicted labels remain non-overlapping within the storage, ensuring the well-definedness and soundness of certification under sample-specific noise. See Algorithm 2 for the overall Cert-SSBD inference procedure, which invokes the storage-update mechanism detailed in Algorithm 3 (Appendix -A).
Although a storage-update-based method is theoretically necessary to guarantee the soundness of certification under sample-specific noise, we did not observe any cases in our experiments where certified regions associated with different predictions overlap. This phenomenon can be attributed to the high dimensionality of image inputs and the moderate magnitude of the optimized noise scales in our evaluated datasets. Nevertheless, in some rare yet realistic scenarios, such overlaps may still occur, particularly when the underlying data distribution includes atypical or ambiguous samples ($e.g.$, label noise or annotation errors, boundary-adjacent inputs with small classification margins, or near-duplicate and highly similar instances). In these cases, our method acts as a conservative and general safeguard: upon the emergence of potential conflicts, it systematically resolves ambiguities by appropriately adjusting the certified regions, thereby preserving the well-definedness and soundness of the resulting certification. Its potential benefits, complete formalization, and additional details are provided in Appendix -C.